OpenStudy (anonymous):

Find the limit of S(n) as n-> infinity s(n) = ∑_(i=1)^n▒(8i-n)/n^2 I keep getting 3 which is not one of the available multiple choice answers. So if someone could double check this I would appreciate it.

OpenStudy (anonymous):

hmm Microsoft word didnt copy correctly The equation is \[s(n) = \sum_{n}^{i=1}(8i-n)/n^2\]

OpenStudy (anonymous):

does it converge?

OpenStudy (anonymous):

I'm not entirely sure what you mean...this is a calc one intro to integration problem...

OpenStudy (anonymous):

I don't think it has a limit. I think it spins off to infinity (and probably beyond)

OpenStudy (anonymous):

and I didnt vet that properly I have my index flipped

OpenStudy (anonymous):

yes, but it doesn't matter too much

OpenStudy (anonymous):

oh ok :(

OpenStudy (anonymous):

what you are really asking (I think) is: \[\sum_{n=1}^{\infty}\sum_{i=1}^{n}{8i-n \over n^2}\]

OpenStudy (anonymous):

your notation is not something I am familiar with but if it means that you are simply substituting \[\infty\] for n then yes

OpenStudy (anonymous):

well that thing does not converge...

OpenStudy (anonymous):

\[(1/n^2)\sum_{i=1}^{n}8i-n\]